OpenStudy (anonymous):

If y = f(x) & y = g(x) intersect at point A on the y-axis, what is the equation of y = f(x) if g(x) = 3x+4 & f'(x) = 2x-3?

OpenStudy (anonymous):

Satellite when your done here may you help me please?

OpenStudy (bahrom7893):

Integrate f'(x) = 2x-3 f(x) = x^2 - 3x +c

OpenStudy (bahrom7893):

Now: since f(x) and g(x) intersect..

OpenStudy (bahrom7893):

they intersect on the y-axis, that means they intersect when g(x) = 0

OpenStudy (bahrom7893):

0 = 3x + 4; 3x = -4 x = -4/3

OpenStudy (bahrom7893):

So now we know that point (-4/3, 0) belongs in there

OpenStudy (bahrom7893):

0 = (-4/3)^2 - 3 (-4/3) +c 0 = 16/9 + 4 + c c = 16/9 - 4 f(x) = x^2 - 3x + (16/9 - 4)

OpenStudy (bahrom7893):

simplify whatever 16/9 -4 is, im doin this in my head.

OpenStudy (anonymous):

Ah thank you very much :)